PrettyEyes:

Roger has a nail that is 12 centimeters long. He measures and records the length of the nail as 15 centimeters. What is the percent error in Roger's measurement? A. 15% B. 17% C. 20% D. 25%

Angle:

what's the difference between the two numbers?

PrettyEyes:

do you mean subtract?

Angle:

yes

PrettyEyes:

ok

PrettyEyes:

3

PrettyEyes:

??

Angle:

awesome! Then we have a formula... \(percent~error=\large \frac{mistake~~difference}{true~~value}\)

Angle:

in this case, 3 goes on top and 12 goes on the bottom

Angle:

\(percent~error=\large \frac{mistake~~difference}{true~~value} = \frac{3}{12} =?\)

PrettyEyes:

hold on ill be back

PrettyEyes:

ok im backwhat is the persent error??

Angle:

\(3 \div 12 \times 100\)

PrettyEyes:

25

Angle:

you got it, that's the answer :)

IsabelFriedmannbells:

great job @PrettyEyes

Angle:

you can also thank people by giving them the "best response" award